OpenStudy (anonymous):

evaluate -4(a+b)-10a/b when a=8 and b=5

OpenStudy (mathteacher1729):

How far have you gotten on this problem?

OpenStudy (anonymous):

i just plugged the numbers in im confused on what to do

OpenStudy (mathteacher1729):

Do you remember the order of operations? Parenthesis Exponents Multiplication & Division Addition & Subtraction (in that order, from left to right) \[-4(8 + 5) - \frac{10(8)}{5}\]

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

ohh imma try it n if i get the answer i will let you know

OpenStudy (mathteacher1729):

Sure! Post the work you did so that if you made any errors, we can point them out. :)

OpenStudy (anonymous):

ok i got -52-16= -68

OpenStudy (mathteacher1729):

Yup, that's correct! :D

OpenStudy (anonymous):

so none of these are the answers 625/8 -53 -15

OpenStudy (mathteacher1729):

Are you sure you copied the problem correctly? In other words, what is the numerator and denominator of that fraction?

OpenStudy (anonymous):

i copied it exactly

OpenStudy (mathteacher1729):

Are you sure it's not this: \[\frac{-4(a+b)-10a}{b}\]

OpenStudy (anonymous):

no

OpenStudy (mathteacher1729):

\[-4(8 + 5) - \frac{10(8)}{5}\] \[=-4(13) - \frac{2(8)}{1}\] \[=-52 - 16\] \[= \fbox{$-68$}\]

OpenStudy (anonymous):

yes thats wat i did

OpenStudy (anonymous):

even if its written the way i wrote it?

OpenStudy (anonymous):

one more question about this one?